a snack is made from granola and nuts...in a 28 oz container there is 1 more ounce of granola than twice the ounces of nuts...
how may ounces of each are in the snack...
i figured out there are 19 oz of granola and 9 oz of nuts, but am not sure how to write the equation....
betsy hoover

Found 2 solutions by ankor@dixie-net.com, ptaylor:
Answer by ankor@dixie-net.com(22740)

a snack is made from granola and nuts...in a 28 oz container there is 1 more ounce of granola than twice the ounces of nuts...
how may ounces of each are in the snack...
:
You're right, here is the algebra way
:
Let x = oz of nuts:
:
It says,"1 more ounce of granola than twice the ounces of nuts." therefore
(2x+1) = oz of granola
:
nuts + granola = 28
x + (2x+1) = 28
3x + 1 = 28
3x = 28 - 1
3x = 27
x = 9 oz of nuts

Answer by ptaylor(2198)

Let x=amount of granola
And y=amount of nuts
Now, we are told that:
amount of granola plus amount of nuts equals 28 oz. So:
x+y=28-----------------------------------eq 1
We are also told that: (remember--one MORE ounce of granola, so we have to take away 1 oz)
x-1=2y---------------------eq 2
In eq 2, add 1 to both sides and we get:
x=2y+1. Substitute this into eq 1:
2y+1+y=28 subtract 1 from both sides
2y+1-1+y=28-1
3y=27 divide both sides by 3
y=9 oz ---------------------------------------------------granola

Substitute y=9 into eq 1 and we have:
x+9=28 subtract 9 from both sides
x=28-9
x=19 oz------------------------------------------------nuts
CK
we know it checks out

Hope this helps----------------------ptaylor